Quote: |
The ~16Hz bounce mode of some optic is showing up in the Yarm error signal.
MC is kind of 'windy' looking, so maybe it's from that? (Yuta's guess).
We need to make sure that the SUS damping and oplev paths both have notches at the correct bounce mode, not the old, old MOS frequency. If that doesn't work, may need to put a resgain in Yarm path.
|
Made the Bounce notch in the BounceRoll filter (ITMY OLPIT, ITMY OLYAW) wider, so it actually spans the peak we see in the error spectra. When we next lock the arm later today, I'll retake this spectra to see if the ETMY oplev fix (Koji, Yuta) and this notch fix both helped. |